# Towards Data-Efficient Pretraining for Atomic Property Prediction

### v1

@@ 3.2 The Chemical Similarity Index (CSI)
Table 1: In-Distribution Evaluation for energy and force targets. We report test MAE when fine-tuning on downstream targets, as detailed in Downstream Datasets (Section 4). The top section represents models pretrained with the large-scale JMP budget, while the lower two sections show results under a limited budget. JMP-S* denotes reproduced results.
@@
$240\times 10^{6}$ Mixed (Temp) JMP-L (GemNet-OC-L) 5.1 1.92 4.75 2.9
JMP-S (GemNet-OC-S) 6.7 2.64 5.71 3.3
JMP-S*(GemNet-OC-S) 6.8 3.21 5.60 3.4
$10\times 10^{6}$ ANI-1x GemNet-OC-S 5.4 2.90 5.13 2.9
@@
$10\times 10^{6}$ Mixed (Balanced) GemNet-OC-S 9.4 3.62 7.02 3.2
Mixed (Temp) 11.0 4.03 7.98 3.6
4 Experiments
Downstream Datasets: For downstream evaluation, we focus on in-distribution (ID) tasks involving energy or force prediction, following the definition in JMP (Shoghi et al., 2023). We discuss out-of-distribution (OOD) tasks in Section 5. Given the large number of pairs for ID evaluation, we focus on the first molecule for force tasks and the energy target for the multi-property dataset QM9 (Wu et al., 2018). The selected targets and their corresponding datasets are: Aspirin in rMD17 (Christensen & Von Lilienfeld, 2020), Ac-Ala3-NHMe in MD22 (Chmiela et al., 2023), solvated amino acids in SPICE (Eastman et al., 2023), and $U_{0}$ in QM9 (Wu et al., 2018).
Baselines: We report the original performance of JMP, where ”JMP-S” and ”JMP-L” correspond to the small and large backbones, respectively. Additionally, we present our reproduced fine-tuning results using the official JMP checkpoints, denoted as ”JMP-S*” and ”JMP-L*”.
For our budgeted evaluation, we present results in two categories: pretraining on a single upstream dataset and pretraining on a mixture of all datasets. For single-dataset experiments, we randomly sample $\mathcal{N}$ instances from the original upstream data. For mixed pretraining, we construct the training set using two different strategies. (1) Balanced Sampling, where an equal number of samples is drawn from each of the four upstream datasets, totaling $\mathcal{N}$ samples; and (2) Temperature-Based Sampling, which preserves the dataset proportions used in the full 120M sample set of JMP (Shoghi et al., 2023).
Evaluation Setup: We pretrain the GemNet-OC-S model (Gasteiger et al., 2022) on each individual upstream dataset, as well as on mixed variants defined in the baselines. For our main experiments, we set a fixed computational budget of $\mathcal{C}=10\times 10^{6}$, achieved by training on $\mathcal{N}=2\times 10^{6}$ samples for $\mathcal{E}=5$ epochs. This budget ensures accessibility and reproducibility, with each pretraining run completing within 1–2 days on an A100 GPU. This represents a 24×24$\times$24 × reduction in computational cost compared to the pretraining budget used in JMP (Shoghi et al., 2023). Additional budget configurations are explored in later sections and the appendix. Each pretrained model is then fine-tuned separately on each downstream task.
4.1 Does CSI Correlate with Better Performance?
In Section 3, we introduced the Chemical Similarity Index (CSI) and presented CSI values. The results indicate that ANI-1x exhibits the highest similarity to all downstream datasets. This finding raises a critical question:
[…]
Table 2: Effect of Computational Budget on Performance. While fixing the number of epochs ($\mathcal{E}$) to 5, we vary the number of training samples from $\mathcal{N}=1\times 10^{6}$ upto $\mathcal{N}=3\times 10^{6}$. Our findings are consistent across budget levels where the upstream dataset with the lowest CSI yields the best downstream performance.
$\mathcal{N}=1\times 10^{6}$ $\mathcal{N}=2\times 10^{6}$ $\mathcal{N}=3\times 10^{6}$
Upstream rMD17 MD22 SPICE QM9 rMD17 MD22 SPICE QM9 rMD17 MD22 SPICE QM9
ANI-1x 5.7 2.91 5.34 2.8 5.4 2.90 5.13 2.9 5.6 2.85 5.37 2.8
Transition-1x 11.7 3.92 7.85 3.3 10.1 3.73 7.55 3.2 10.2 3.64 7.65 3.2
OC20 14.8 4.67 8.92 4.7 14.6 4.53 8.74 4.8 16.1 4.61 10.44 5.3
OC22 18.1 5.44 11.22 5.5 16.0 5.20 10.73 5.7 17.4 5.15 11.00 5.8
Table 3: Effect of Changing the Backbone Size. We analyze the impact of using a larger variant of GemNet-OC and find that, irrespective of backbone size, relevance-based upstream dataset selection consistently outperforms costly large-scale dataset mixing.
$\mathbf{\mathcal{C}}$ Upstream Data Backbone rMD17 MD22 SPICE QM9
(meV/Å) (meV/Å) (meV/Å) (meV)
$240\times 10^{6}$ Mixed (Temp) JMP-L (GemNet-OC-L) 5.1 1.92 4.75 2.9
JMP-L* (GemNet-OC-L) 5.3 2.59 4.91 3.0
$10\times 10^{6}$ ANI-1x GemNet-OC-L 4.8 2.54 5.24 2.6
Transition-1x 9.7 3.56 7.42 3.0
OC20 13.8 3.90 9.24 4.6
OC22 12.0 4.14 10.43 4.0
Table 1 summarizes the downstream performance of models pretrained on different datasets in an in-distribution setting. Notably, the ANI-1x model consistently outperforms all other individual datasets and even mixed variants. For instance, on both rMD17, SPICE and QM9 datasets, ANI-1x achieves an MAE of 5.4, 5.13 and 2.9, compared to 6.7, 5.71 and 3.3 for JMP-S. Remarkably, this performance is achieved with 24×24$\times$24 × less computational budget than JMP-S. Additionally, pretraining on alternative upstream tasks results in weaker performance compared to ANI-1x, correlating with the trends observed in our CSI values.
Furthermore, temperature-based mixing, which follows the JMP formulation and prioritizes high-CSI datasets (OC20 and OC22), performs worse. In contrast, balanced mixing improves slightly by incorporating more samples from the more relevant ANI-1x and Transition-1x but remains inferior to individual pretraining on ANI-1x. These results indicate that, under a limited budget, mixing upstream datasets with varying CSI values is suboptimal and demands significantly more computational resources for convergence.
Takeaway. Our experiments highlight three key insights for in-distribution downstream tasks: (1) High-quality, task-relevant upstream datasets like ANI-1x outperform larger, mixed datasets, when included within the mix. (2) Mixing upstream datasets can match the benefits of highly relevant pretraining but requires significantly more compute and training time. (3) CSI effectively predicts downstream performance, as lower CSI values (e.g., ANI-1x) consistently correlate with better results.
4.2 What is the Effect of Computational Budget?
[…]
Do our findings about dataset relevance in terms of CSI hold across different budget levels?
Table 2 presents the downstream performance across additional pretraining budgets of $1M$, and $3M$ samples (5 epochs each) compared to $2M$. ANI-1x remains superior to other datasets across all budgets, aligning with its low CSI. Increasing the budget from $1M$ to $2M$ samples generally improves performance; however, further scaling to $3M$ samples results in diminishing returns on rMD17, SPICE, and QM9, with only a slight improvement on MD22. This occurs despite the pretraining loss for $3M$ being better than $2M$, highlighting that improved pretraining performance does not always translate to better downstream results.
4.3 What is the Effect of Changing the Architecture Size?
[…]
Does the correlation between CSI and downstream performance hold across different architecture sizes?
Table 3 reports the downstream performance when using the large variant, GemNet-OC-L, as the backbone. Consistent with the results on the small backbone, models pretrained on ANI-1x achieve the best performance, correlating well with the CSI values. Notably, we achieve state-of-the-art results with an MAE of 4.8 on Aspirin in rMD17 and 2.6 on $U_{0}$ in QM9, even though our reproduced JMP-L baseline results fall short of those reported in the original JMP paper.
Takeaway. Our findings hold across backbone sizes, showing the potential of relevance-based upstream dataset selection over large-scale mixing for improved performance.
Figure 4: Impact of Adding Less Relevant Pretraining Data. Adding $1M$ OC22 samples to a $2M$-sample ANI-1x baseline worsens downstream performance despite a larger pretraining budget. This highlights the importance of dataset relevance and the CSI metric for effective pretraining.
Table 4: OOD Task Performance Across Upstream Sources. We compare the CSI-predicted best upstream sources with actual downstream performance on OOD tasks (QMOF, MatBench, and QM9’s $\Delta_{\epsilon}$). While CSI aligns well with QM9’s OOD label, it mispredicts the best source for MatBench. Mixed pretraining generally improves performance, highlighting the benefits of diverse upstream sources for OOD generalization.
$\mathbf{\mathcal{C}}$ Upstream Data Backbone QM9 [$\Delta_{\epsilon}$] QMOF MatBench [fold0 / mean]
($meV$) (eV) ($cm^{-1}$)
$240\times 10^{6}$ Mixed (Temp) JMP-S (GemNet-OC-S) 23.1 0.18 26.60 / 22.77
JMP-S* (GemNet-OC-S) 24.0 0.19 24.77 / 21.48
$10\times 10^{6}$ ANI-1x GemNet-OC-S 24.5 0.22 30.09 / 29.60
Transition-1x 25.3 0.22 52.22 / 38.56
OC20 30.8 0.22 37.52 / 30.88
OC22 35.6 0.22 32.78 / 27.55
$10\times 10^{6}$ Mixed (Balanced) GemNet-OC-S 27.3 0.21 26.11 / 24.87
Mixed (Temp) 27.9 0.21 26.63 / 25.61
4.4 Is More Data Always Better?
The common pretraining paradigm assumes that larger and more diverse datasets lead to better generalization. Here, we challenge this assumption from another perspective:
[…]
Does adding pretraining data from less relevant sources improve or degrade downstream performance?
To evaluate this, we pretrain on a mixture of $2M$ samples from the best CSI dataset (ANI-1x) and $1M$ from OC22, a non-aligned upstream dataset. We then compare this pretraining checkpoint with the $2M$-sample baseline in Figure 4. Surprisingly, adding OC22 degraded downstream performance despite the increased pretraining budget.
Takeaway. Including additional pretraining data from less relevant sources can harm downstream performance. Our findings highlight the CSI metric as a practical tool for designing effective pretraining datasets.
5 Beyond In-Distribution
Recall that our pretraining process is conducted on upstream tasks involving molecules and catalysts, with energy and force as targets. For downstream tasks with different labels (e.g., band gap in QMOF) or from distinct chemical domains such as materials (e.g., MatBench and QMOF), we classify these as out-of-distribution (OOD). While our main results focused on ID evaluation, here we explore our metric’s applicability to OOD tasks. Specifically, we examine three cases: the Band Gap property from QMOF (Rosen et al., 2021), Phonons (the first non-energy target in JMP tables) from MatBench (Dunn et al., 2020), and $\Delta_{\epsilon}$ from QM9, explicitly categorized as OOD in the JMP paper.
In Figure 5, we present the CSI values for OOD domains, where the OOD label ($\Delta_{\epsilon}$) for QM9 follows the same values as in Figure 3. We observe that QMOF exhibits a pattern similar to other ID domains shown in Figure 3. However, MatBench displays a distinct pattern, showing strong correlation with OC20 and OC22, followed by ANI-1x and Transition-1x. Next, we analyze the correlation between CSI and downstream performance under OOD evaluation.
Table 4 shows that $\Delta_{\epsilon}$ in QM9 aligns with the CSI pattern, similar to ID evaluation, suggesting that CSI is effective for OOD in the label space. In QMOF, the different upstream sources achieve similar performance which lags behind the full pretraining by JMP. For MatBench (evaluated over 5 folds), OC22 achieves the best mean performance while OC20 lags behind, despite our metric predicting both to be equally suitable. Additionally, for both QMOF and MatBench, mixed pretraining variants generalize better than individual sources. This suggests that when the downstream domain differs from all upstream sources, mixing diverse upstream domains provides the best performance.
While our CSI metric holds well for ID evaluation, further research is needed to improve OOD generalization across different chemical domains. One direction is to explore additional upstream sources to better capture domain variations. Another potential extension is leveraging foundation models trained beyond energy and forces, which could enhance feature extraction and improve similarity assessments.
6 Conclusion
This paper challenges the prevailing trend of scaling data and computational resources in atomic property prediction by demonstrating that strategic data selection based on relevance can achieve comparable or superior performance with significantly fewer resources. We introduce the Chemical Similarity Index (CSI), a simple metric that quantifies upstream/downstream datasets alignment, enabling the selection of high-quality, task-relevant pretraining data. Our experiments reveal that smaller, focused datasets often outperform larger, mixed ones, and that indiscriminately adding data can degrade performance when relevance is low. These findings highlight that quality often outweighs quantity in pretraining, offering a more efficient and sustainable path forward for molecular machine learning.
Acknowledgements
The authors thank Mohamed Elhoseiny for his feedback. Yasir Ghunaim was supported by Saudi Aramco. The research reported in this publication was supported by funding from King Abdullah University of Science and Technology (KAUST) - Center of Excellence for Generative AI, under award number 5940 and SDAIA-KAUST Center of Excellence in Data Science, Artificial Intelligence (SDAIA-KAUST AI).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A More Epochs or More Data?
In this section, we explore the trade-off between increasing the number of training epochs and expanding the dataset size under a fixed computational budget. Specifically, we aim to answer the following question:
[…]
Interestingly, increasing the number of epochs from 5 to 10 while reducing the dataset size from $2M$ to $1M$ does not significantly degrade performance for ANI-1x. This suggests that for highly relevant datasets like ANI-1x, training on fewer samples for more epochs can be as effective as training on more samples for fewer epochs. However, for less relevant datasets like OC20 and OC22, increasing the number of epochs does not compensate for the reduced dataset size, as their performance degrades significantly.
Conclusion. Our findings indicate that the choice between more epochs or more data depends on the relevance of the dataset to the downstream task. For highly relevant datasets like ANI-1x, training on fewer samples for more epochs can yield comparable or even superior performance, while for less relevant datasets, increasing the dataset size is more beneficial. This further underscores the importance of dataset quality and relevance, as quantified by CSI, in determining the optimal pretraining strategy.
$\mathbf{\mathcal{C}}$ $\mathbf{\mathcal{N}}$${\mathcal{E}}$ Upstream Data Backbone rMD17 MD22 SPICE QM9
(meV/Å) (meV/Å) (meV/Å) (meV)
$10\times 10^{6}$ $2\times 10^{6}$ $5$ ANI-1x GemNet-OC-S 5.4 2.90 5.13 2.9
Transition-1x 10.1 3.73 7.55 3.2
OC20 14.6 4.53 8.74 4.8
OC22 16.0 5.20 10.73 5.7
$10\times 10^{6}$ $1\times 10^{6}$ $10$ ANI1x GemNet-OC-S 5.4 2.88 5.04 2.9

### v2

@@ 4.1 Does CSI Correlate with Better Performance?
Table 1: In-Distribution Evaluation for Energy and Force Targets. We report test MAE when fine-tuning on downstream targets, as detailed in Downstream Datasets (Section 4). The top section represents models pretrained with the large-scale JMP budget, while the lower two sections show results under a limited budget. JMP-S* denotes reproduced results. We report the average and standard deviation over three random seeds.
$\mathbf{\mathcal{C}}$ Upstream Data Backbone rMD17 MD22 SPICE QM9 OMat24
(meV/Å) (meV/Å) (meV/Å) (meV) (meV/Å)
240M Joint (Temperature) JMP-L (GemNet-OC-L) 5.1 1.92 4.75 2.9 -
JMP-S (GemNet-OC-S) 6.7 2.64 5.71 3.3 -
JMP-S*(GemNet-OC-S) 6.8 3.21 5.60 3.4 84.4
10M ANI-1x GemNet-OC-S 5.2 $\pm$ 0.2 2.92 $\pm$ 0.03 5.39 $\pm$ 0.26 3.1 $\pm$ 0.1 98.1 $\pm$ 0.5
Transition-1x 9.8 $\pm$ 0.5 3.70 $\pm$ 0.02 7.70 $\pm$ 0.31 3.6 $\pm$ 0.3 100.8 $\pm$ 0.1
OC20 13.8 $\pm$ 0.7 5.10 $\pm$ 0.57 9.67 $\pm$ 0.73 5.3 $\pm$ 0.4 87.4 $\pm$ 0.1
OC22 15.3 $\pm$ 0.7 5.34 $\pm$ 0.18 10.62 $\pm$ 0.42 5.7 $\pm$ 0.1 92.5 $\pm$ 0.1
10M Joint (Balanced) GemNet-OC-S 9.1 $\pm$ 1.8 3.73 $\pm$ 0.13 6.99 $\pm$ 0.33 3.5 $\pm$ 0.2 90.3 $\pm$ 0.1
Joint (Temperature) 11.0 $\pm$ 1.5 4.41 $\pm$ 0.59 8.44 $\pm$ 0.63 4.0 $\pm$ 0.4 88.4 $\pm$ 0.1
Furthermore, joint pretraining, whether balanced or temperature-based, generally performs worse than individual pretraining on the most CSI-aligned dataset for each task when both are trained under the same limited-budget setting. Following the JMP formulation, temperature-based sampling allocates more samples to OC20 and OC22 than balanced sampling. On OMat24, where OC20 and OC22 are most aligned according to CSI values, temperature-based sampling outperforms balanced sampling. However, temperature-based sampling degrades performance on other tasks where OC20 and OC22 are less aligned with the downstream dataset. These results suggest that, under a limited budget, mixing upstream datasets with varying CSI values is suboptimal and requires significantly larger pretraining budgets to achieve competitive performance.
Takeaway. Our experiments reveal three key insights for in-distribution downstream tasks: (1) Task-aligned upstream datasets outperform larger joint datasets. (2) Joint pretraining can match the benefits of pretraining on a highly task-aligned dataset, but doing so requires substantially larger pretraining budgets. (3) CSI effectively predicts downstream performance, as lower CSI values consistently correlate with better results.
4.2 What is the Effect of Computational Budget?
Do our findings about dataset alignment in terms of CSI hold across different budget levels?
Figure 4 shows downstream MAE across pretraining budgets of 0.5M, 1M, 2M, and 3M samples (each trained for 5 epochs). We observe that increasing the budget for low-CSI datasets (i.e., ANI-1x and Transition-1x) beyond 2M leads to diminishing returns on rMD17 and SPICE. In contrast, increasing the pretraining budget for high-CSI datasets (i.e., OC20 and OC22) often degrades downstream performance more substantially, particularly on rMD17, SPICE, and QM9. These results indicate that allocating more compute to misaligned upstream tasks can harm downstream generalization.
Figure 4: Effect of Computational Budget on Performance. While fixing the number of epochs ($\mathcal{E}$) to 5, we vary the number of training samples across $\mathcal{N}=0.5$M, 11M, 22M, and 33M. Our findings are consistent across budget levels where the upstream dataset with the lowest CSI yields the best downstream performance.
Table 2: Effect of Changing the Backbone Size. We analyze the impact of using a larger variant of GemNet-OC and find that, irrespective of backbone size, relevance-based upstream dataset selection consistently outperforms costly large-scale joint pretraining.
@@ 4.2 What is the Effect of Computational Budget?
240M Joint (Temperature) JMP-L (GemNet-OC-L) 5.1 1.92 4.75 2.9
JMP-L* (GemNet-OC-L) 5.3 2.59 4.91 3.0
10M ANI-1x GemNet-OC-L 4.8 2.54 5.24 2.6
Transition-1x 9.7 3.56 7.42 3.0
OC20 13.8 3.90 9.24 4.6
OC22 12.0 4.14 10.43 4.0
10M Joint (Balanced) GemNet-OC-L 8.9 3.39 7.57 3.2
Joint (Temperature) 10.8 4.15 9.83 4.4
4.3 What is the Effect of Changing the Backbone Size?
Does the correlation between CSI and downstream performance hold across different backbone sizes?
Table 2 reports the downstream performance using the large variant, GemNet-OC-L, as the backbone. We also include our best attempt at reproducing the baseline results using JMP-L pretraining (denoted as "JMP-L*").
Consistent with the results on the small backbone, models pretrained on ANI-1x achieve the best performance across all downstream tasks, aligning with its low CSI values. Notably, using a small computation budget of $\mathcal{C}=10$M (i.e., 2M samples over 5 epochs), ANI-1x outperforms JMP-L, which was pretrained with $\mathcal{C}=240$M on a joint upstream dataset. We obtain state-of-the-art results with an MAE of 4.8 on Aspirin (rMD17) and 2.6 on $U_{0}$ (QM9), demonstrating that strong dataset alignment can outweigh large-scale pretraining even with increased model capacity. While larger backbones improve overall performance, the gap between aligned and misaligned upstream datasets persists. High-CSI datasets like OC20 and OC22 still underperform, reaffirming the importance of dataset alignment.
Takeaway. Our findings hold across backbone sizes: scaling up the model does not change the relative utility of upstream datasets. Alignment-based upstream dataset selection outperforms large-scale dataset mixing, even under high-capacity settings and at significantly lower computational budgets.
4.4 What is the Effect of Changing the Backbone Type?
In the previous experiments, we focused on the GemNet-OC backbone to enable a direct and fair comparison with the JMP baseline. We now examine whether our CSI-based alignment analysis extends beyond graph neural networks to a different architectural paradigm. Specifically, we ask:
Does the relationship between CSI and downstream performance persist across different backbone types?
To answer this question, we evaluate EquiformerV2 (Liao et al., 2023), a transformer-based equivariant backbone. We pretrain EquiformerV2 on the same set of upstream datasets using the identical pretraining and fine-tuning protocol as in our main experiments. We adopt the same pretraining budget of 2M samples over 5 epochs for both the small (31M parameters) and large (87M parameters) EquiformerV2 variants. Pretraining EquiformerV2 is notably more expensive than GemNet-OC. Even under this limited 10M budget, the large variant requires up to 17 A100 GPU-days. We report the results in Table 3.
Across both EquiformerV2 model sizes, performance trends remain consistent with our GemNet-OC findings. The datasets with the lowest CSI lead to the strongest downstream performance, with ANI-1x being the top-performer in nearly all cases. While Transition-1x shows a marginal advantage on QM9 for the 31M model, ANI-1x achieves the best results as capacity increases to the 87M variant. This confirms that the benefits of alignment reflect a general property of data relevance that persists across architectural paradigms.
Table 3: Effect of Changing the Backbone Type. Downstream MAE after fine-tuning EquiformerV2 (EqV2) pretrained on different upstream datasets. Despite the architectural shift from graph neural networks to transformer-based models, upstream dataset alignment remains a strong predictor of downstream performance. Numbers in parentheses indicate parameter count.
$\mathbf{\mathcal{C}}$ Pretraining Time Upstream Data Backbone rMD17 MD22 SPICE QM9
(A100 GPU-days) (meV/Å) (meV/Å) (meV/Å) (meV)
10M 1.68 ANI-1x EqV2 (31M) 9.7 3.23 8.88 6.5
1.60 Transition-1x 11.0 3.37 10.16 6.4
6.78 OC20 12.5 3.58 10.54 7.0
7.72 OC22 12.6 3.72 11.10 6.7
10M 6.69 ANI-1x EqV2 (87M) 9.0 2.90 6.24 3.5
6.72 Transition-1x 10.2 3.05 7.80 3.6
16.52 OC20 11.4 3.25 8.10 4.0
17.16 OC22 11.9 3.28 8.60 4.3
Takeaway. CSI-based alignment reliably predicts downstream performance across both backbone sizes and backbone types, highlighting the central role of dataset alignment in determining pretraining performance.
4.5 Is More Diverse Data Always Better?
A common assumption in pretraining is that larger and more diverse datasets lead to better generalization. This intuition motivates the JMP framework, where a large-scale pretraining budget of $\mathcal{C}=240$M led to strong downstream results. However, it remains unclear whether this benefit comes from the size, the diversity, or the alignment of the data with the downstream task. Here, we revisit this assumption through a targeted experiment:
Does increasing data diversity by adding less aligned sources improve or harm downstream performance?
To test this, we compare two settings: (1) pretraining on $\mathcal{N}=2$M unique samples from ANI-1x, the most CSI-aligned dataset, and (2) pretraining on a mixture of 2M ANI-1x samples and 1M OC22 samples (i.e., $\mathcal{N}=3$M), both trained for 5 epochs. As shown in Figure 5, simply adding OC22 results in worse downstream performance across all four tasks, despite the increase in data volume. This indicates that adding less aligned data may interfere with the knowledge transfer gained from aligned pretraining sources.
Figure 5: Impact of Adding Less Aligned Pretraining Data. Adding $1M$ OC22 samples to a $2M$-sample ANI-1x baseline worsens downstream performance despite a larger pretraining budget. This highlights the importance of dataset alignment and the value of the CSI metric for effective pretraining.
Takeaway. Our results challenge the intuitive strategy of adding diversity to pretraining datasets without considering alignment. CSI provides a practical signal for curating upstream data that supports better generalization, especially under constrained budgets.
4.6 Is Alignment Important in Self-Supervised Pretraining?
Self-supervised learning has long been central to 3D molecular representation learning, initially driven by the scarcity of large-scale labeled data (Rong et al., 2020; Liu et al., 2021; Jiao et al., 2023; Chen et al., 2021; Zhou et al., 2022a; Ji et al., 2024). While recent frameworks such as JMP (Shoghi et al., 2023) emphasize supervised pretraining on energy and forces, modeling structural dependencies directly from molecular geometry remains a dominant paradigm. This raises a critical question regarding the role of data selection:
Does dataset alignment matter for self-supervised pretraining?
To answer this, we extend our analysis to the self-supervised setting and adopt the DeNS framework (Liao et al., 2024). DeNS follows the standard denoising paradigm in 3D molecular learning, where models are trained to recover the original 3D atomic positions from noisy inputs. Unlike traditional self-supervised denoising objectives designed for equilibrium geometries (Rong et al., 2020; Zhou et al., 2022a), DeNS explicitly targets non-equilibrium molecular and material configurations. This makes it well suited to the non-equilibrium datasets considered in this work. We pretrain the EquiformerV2 (87M) backbone with the DeNS objective on each upstream dataset using a fixed budget of 10M training instances (2M samples over 5 epochs).
Table 4: Effect of Changing the Pretraining Objective. We report the test MAE for the EquiformerV2 (EqV2) backbone, with 87M parameters, pretrained with the DeNS self-supervised denoising objective. Consistent with our supervised results, pretraining on the most CSI-aligned dataset (i.e., ANI-1x) achieves superior fine-tuning performance across all downstream molecular tasks.
$\mathbf{\mathcal{C}}$ Upstream Data Backbone rMD17 MD22 SPICE QM9
(meV/Å) (meV/Å) (meV/Å) (meV)
10M ANI-1x EqV2 (87M) + DeNS 4.7 2.78 5.95 4.1
Transition-1x 7.3 2.95 8.71 4.2
OC20 8.1 3.32 7.34 5.0
OC22 10.4 3.58 9.39 4.9
[…]
Takeaway. Dataset alignment is objective-agnostic. Even when the pretraining objective is changed to model structural dependencies in a self-supervised manner, CSI-aligned datasets consistently achieve superior performance.
Table 5: OOD Task Performance Across Upstream Sources. We compare the CSI-predicted best upstream sources with actual downstream performance on OOD tasks (QMOF, MatBench, and QM9’s $\Delta_{\epsilon}$). While CSI aligns well with QM9’s OOD label, it mispredicts the best source for MatBench. Joint pretraining generally improves performance, highlighting the benefits of diverse upstream sources for OOD generalization.
$\mathbf{\mathcal{C}}$ Upstream Data Backbone QM9 $\Delta_{\epsilon}$] QMOF MatBench [fold0 / mean]
(meVmeV) (eV) ($cm^{-1}$)
240M Joint (Temperature) JMP-S (GemNet-OC-S) 23.1 0.18 26.60 / 22.77
JMP-S* (GemNet-OC-S) 24.0 0.19 24.77 / 21.48
10M ANI-1x GemNet-OC-S 24.5 0.22 30.09 / 29.60
Transition-1x 25.3 0.22 52.22 / 38.56
OC20 30.8 0.22 37.52 / 30.88
OC22 35.6 0.22 32.78 / 27.55
10M Joint (Balanced) GemNet-OC-S 27.3 0.21 26.11 / 24.87
Joint (Temperature) 27.9 0.21 26.63 / 25.61
5 Beyond In-Distribution
Recall that our pretraining process is conducted on upstream tasks involving molecules and catalysts, with energy and force as targets. We classify as in-distribution (ID) any downstream task that uses the same labels (energy and/or forces), and as out-of-distribution (OOD) those that involve different target properties, such as band gap (e.g., QMOF) or phonon properties (e.g., MatBench). While our main results focused on ID evaluation, here we explore our metric’s applicability to OOD tasks. Specifically, we examine three cases: the Band Gap property from QMOF (Rosen et al., 2021), Phonons from MatBench (Dunn et al., 2020), and $\Delta_{\epsilon}$ from QM9.
In Figure 6, we present the CSI values for OOD domains; for QM9, the OOD label ($\Delta_{\epsilon}$) uses the same dataset-level CSI values as in Figure 3. We observe that QMOF exhibits a pattern similar to other ID domains shown in Figure 3. However, MatBench displays a distinct pattern, showing strong correlation with OC20 and OC22, followed by ANI-1x and Transition-1x. Next, we analyze the correlation between CSI and downstream performance under OOD evaluation.
Table 5 shows that $\Delta_{\epsilon}$ in QM9 aligns with the CSI pattern, similar to ID evaluation, suggesting that CSI remains effective when the downstream domain (i.e., molecules, as in QM9) is similar to the upstream tasks, even if the label type is OOD. In QMOF, the different upstream sources achieve similar performance, which lags behind the full pretraining by JMP. For MatBench (evaluated over 5 folds), OC22 achieves the best mean performance while OC20 lags behind, despite our metric predicting both to be equally suitable. Additionally, for both QMOF and MatBench, joint pretraining variants generalize better than individual sources. This suggests that when both the downstream domain and the label type differ from all upstream sources, mixing diverse upstream domains provides the best performance.
While CSI reliably guides dataset selection for in-distribution tasks, its effectiveness in OOD scenarios reveals important limitations of dataset-level alignment. This may stem from the limited diversity of the backbone used for feature extraction, which was pretrained only on energy and force targets. Future work could explore using backbones pretrained on broader sets of chemical properties or incorporating more diverse upstream domains to better capture variation across OOD tasks. Another promising direction is to leverage foundation models trained on multi-modal or multi-objective tasks, which may offer more transferable representations for similarity assessment across varied downstream domains.
6 Conclusion
This paper challenges the prevailing trend of scaling data and computational resources in atomic property prediction by demonstrating that strategic data selection based on dataset alignment can achieve comparable or superior performance with significantly fewer resources. We introduce the Chemical Similarity Index (CSI), a simple metric that quantifies the alignment between upstream pretraining datasets and downstream tasks, enabling the selection of high-quality, task-aligned pretraining data. Our experiments reveal that smaller, focused datasets often outperform larger, mixed ones, and that indiscriminately adding data can degrade performance when relevance is low. These findings highlight that alignment, rather than scale alone, is the key to effective pretraining, and they point toward a more principled, efficient, and sustainable direction for future research in atomic property prediction.
Acknowledgment
The authors thank Mohamed Elhoseiny, Firas Laakom, and Abdelrahman Eldesokey for their valuable feedback. Yasir Ghunaim was supported by Saudi Aramco. This work is supported by the King Abdullah University of Science and Technology (KAUST) Center of Excellence for Generative AI under award number 5940 and SDAIA-KAUST Center of Excellence in Data Science, Artificial Intelligence (SDAIA-KAUST AI). The computational resources are provided by IBEX, which is managed by the Supercomputing Core Laboratory at KAUST.
Appendix A More Epochs or More Data?
To extend the findings presented in the main paper, we explore the trade-off between increasing the number of training epochs and expanding the dataset size under a fixed computational budget. Specifically, we aim to answer the following question:
Interestingly, increasing the number of epochs from 5 to 10 while reducing the dataset size from 2M to 1M does not significantly degrade performance for ANI-1x. This suggests that for highly aligned datasets like ANI-1x, training on fewer samples for more epochs can be as effective as training on more samples for fewer epochs. In contrast, for less aligned datasets such as OC20 and OC22, increasing the number of epochs only partially compensates for the reduced dataset size, as some tasks show similar performance while others experience noticeable degradation.
Takeaway. Our findings indicate that the trade-off between more epochs and more data depends on the alignment of the pretraining dataset with the downstream task. For highly aligned datasets like ANI-1x, training on fewer samples for more epochs can yield comparable performance. In contrast, for less aligned datasets, increasing the dataset size tends to be more beneficial. These results further show the importance of dataset quality and alignment, as quantified by CSI, in determining an effective pretraining strategy.
$\mathbf{\mathcal{C}}$ $\mathbf{\mathcal{N}}$ $\mathbf{\mathcal{E}}$ Upstream Data Backbone rMD17 MD22 SPICE QM9
(meV/Å) (meV/Å) (meV/Å) (meV)
10M 2M 5 ANI-1x GemNet-OC-S 5.4 2.90 5.13 2.9
@@ Appendix A More Epochs or More Data?
10M 1M 10 ANI1x GemNet-OC-S 5.4 2.88 5.04 2.9